# The Three Hundred Project: deducing the stellar splashback structure of galaxy clusters from their orbiting profiles

### v1

@@ 4.2 Recovering $r_{\rm t}$ from observable profiles
We calculate the projected density profiles using the same radial binning as we used for the 3D profiles. The angular binning is also done with the same number of bins, but this time in equal angle rather than equal solid angle. We project the entire cluster region along the $x$, $y$ and $z$ axes, and take the mean of the resulting profiles. While choosing different windows along the line of sight will produce distinct profiles, we mainly concern ourselves with the recoverability of the orbiting parameters rather than their actual values in projection. The calculation of the density slope profiles is the same as for the 3D case. For clarity, we refer to the projected density with the symbol $\Sigma$ and other quantities with upper case, e.g. $R_{\rm t}$ and $B$ for the truncation radius and steepness.
In projection, the shape of the density profile is modified. For the stars, the splashback region and BCG-stellar halo transition are smoothed out. This can be seen in Figure 6, which compares the 3D and projected stellar profile of one of the clusters in the GIZMO_7k run. Despite this, the profiles are still well-fit by the functions 1 - 3, generalizing earlier work on the similarity between 3D and projected total profiles (e.g. Merritt et al. 2005). It is reasonable to expect that the parameters of the 3D and projected profiles should be related. Figure 7 compares the truncation radius and truncation steepness from fitting the 3D and projected orbiting profiles (not the total profiles), for the dark matter (left) and stars (right). The projected $R_{\rm t}$ is slightly smaller than the 3D $r_{\rm t}$ by a factor of $\sim 0.92$ for the dark matter and $0.94$ for the stars (black dashed lines). This is similar to the factor of $0.9$ found by Deason et al. (2021) when comparing the 3D and projected steepest slope radii. The projected $B$ of the dark matter is also smaller than its 3D counterpart by a factor of $\sim 0.87$. This may also be the case for the stars, but the scatter is significantly larger.
@@
Figure 8: Comparison between $r_{\rm t}$ obtained by fitting the orbiting profile (component) and the total profile. The points are colored by accretion rate (Equation 4). The cyan squares are from the $\Gamma$-binned median profiles.
Figure 9: $R_{\rm steep}$ and $R_{\rm t}$ as a function of accretion rate (Equation 4). The red squares are from fitting the $\Gamma$-binned median profiles. The black dashed line in the top row is the relation from More et al. (2015), rescaled using the factor of $r_{\rm steep}/R_{\rm steep}=0.9$ found by Deason et al. (2021). The black dashed in the bottom row is the same curve further rescaled using the best-fit $r_{\rm t}$-$r_{\rm steep}$ relation shown in Figure 4. The black dotted line is the relationship from Diemer (2025), rescaled by the slope of the best-fit line shown in Figure 7.
Figure 10: Best-fit parameters for the 3D infalling profiles. The bottom-left (red panels) and top-right (purple panels) show the correlations between each parameter for the dark matter and stars, respectively. The gray diagonal panels compare the parameters between the stars and dark matter. The green and red squares are the parameters of the mean and median (respectively) $\Gamma$-binned profiles.
@@
The $R_{\rm steep}$-$\Gamma$ plots feature a small group of points that lie below the relation at low $\Gamma$. These are cases where the $2^{\rm nd}$ caustic—an additional, usually shallower caustic at smaller radius—actually becomes steeper than the $1^{\rm st}$ caustic when viewed in projection. The measurement of $R_{\rm t}$ seems to be unaffected by this, suggesting the fitting method may have an advantage in these cases.
@@ 4.3 The infalling profile
Figure 10 shows the distribution of the parameters from fitting the 3D infalling profiles. The bottom-left (red panels) and top-right (purple panels) show the correlations between each parameter for the dark matter and stars, respectively. The gray diagonal panels compare the parameters between the stars and dark matter. The green and red squares are the parameters of the mean and median (respectively) $\Gamma$-binned profiles.
@@ 4.4 The inner profile
In our model, the orbiting density profile of the stars is composed of a Sérsic BCG and a stellar “halo” fit with the truncated Einasto profile of Diemer (2025) (see Section 3.3). As with the unmodified Einasto profile, the shape of the inner profile is controlled by the inner slope exponent $\alpha$ and the scale radius $r_{\rm s}$—the radius at which the logarithmic slope equals 2-2. Figure 12 compares $r_{\rm s}$ and $\alpha$ between the stars and dark matter. Unlike for the outer profile parameters, the stellar inner profile parameters do not track the dark matter well. Though there is a population with $r_{\rm s}$ along the one-to-one line, the stellar $r_{\rm s}$ tends to underestimate that of the dark matter, while $\alpha$ has significant scatter around the one-to-one line.
Figure 12: Comparison between $r_{\rm s}$ and $\alpha$ obtained from fitting the dark matter and stellar orbiting profiles. The points are colored by accretion rate (Equation 4). The cyan squares show the results from fitting the $\Gamma$-binned median profiles.
Figure 13: Comparison between stellar and dark matter $\rho r^{2}$ profiles of the nine most massive clusters in the GIZMO_7k run. The stellar profiles have been rescaled by a fixed number $k_{\rm SHM}=138$. The sub-panels show the derivative of the profiles. The vertical gray lines mark the location of the peak of the dark matter profile.
The reason for this disagreement is the freedom in the inner profile allowed by the BCG term. The BCG is typically and order of magnitude more dense than the stellar halo (see Figures 2 and 6), meaning that, in log space, even a small change to the BCG fit can lead to a large change in the shape of the inner stellar halo profile.
To see if the disagreement between the stellar and dark matter $r_{\rm s}$ suggested by Figure 12 is real or merely an artifact of the fitting, we measure $r_{\rm s}$ non-parametrically by plotting the $\rho\,r^{2}$ profile, which peaks when the logarithmic slope of $\rho$ is 2-2. Figure 13 shows the $\rho\,r^{2}$ profiles of nine of the most massive clusters in the GIZMO_7k run, along with their logarithmic derivatives. The stellar profiles have been rescaled by a fixed factor of $k_{\rm SHM}=138$. The peaks in the stellar profile are more difficult to discern than those of the dark matter due to their proximity to the central density enhancement—nevertheless, in most cases they are very close to those of the dark matter, suggesting that the $r_{\rm s}$ of the stars should indeed coincide with that of the dark matter.
@@ 5.1 Comparison to Previous Simulations
The coincidence between the $r_{\rm t}$ means that the stellar $r_{\rm t}-\Gamma$ relation has the same form as that of the dark matter. This makes it a complimentary estimator of the accretion rate in clusters—assuming that $r_{\rm t}$ can be measured accurately. However, in observations, information of the orbiting and infalling components is inaccessible, and $r_{\rm t}$ must instead be estimated by fitting the total profile in projection. Given the freedom allowed by the full nine-parameter model of the total profile, the orbiting and infalling components may not always be accurately recovered.
When fitting the total profile, we find that the projected truncation radius, $R_{\rm t}$, has a standard deviation from the “true” $R_{\rm t}$ (measured from the orbiting profile) of $\sim 0.3\,R_{200{\rm m}}$ for both the dark matter and stars. There may also be a slight $R_{\rm t}$-dependent systematic difference in the case of the stars, which would produce a subtle change in the $R_{\rm t}$-$\Gamma$ relation, but this is unlikely to be discernable. While the exponential relationship can still be clearly seen, the scatter is larger than that of the $R_{\rm steep}$-$\Gamma$ relationship and so it is likely to be a poorer estimator of the accretion rate in observations.
@@ 5.2 Comparison to Observations
Gravitational lensing, Sunyaev–Zel’dovich (SZ) observations, and measurements of the distribution of cluster galaxies in projection provide complementary approaches to identifying the splashback boundary in galaxy clusters, each offering independent insight into the halo’s outer structure. Weak lensing directly probes the total matter distribution and has revealed a steepening in the density profile consistent with theoretical predictions of the splashback feature (e.g. Chang et al., 2018b; Contigiani et al., 2019). However, lensing measurements are inherently noisy at large radii due to shape noise and projection effects, and often require stacking to obtain robust profiles.
Chang et al. (2018b) conducted one of the earliest joint analyses of galaxy density and weak lensing profiles using DES Year 1 data, detecting a splashback-like steepening in both tracers. They measured a splashback radius of $r_{\mathrm{sp}}=1.34\pm 0.21\,h^{-1}\,\mathrm{Mpc}$, consistent with predictions from simulations and earlier SDSS-based studies. This dual detection reinforced the interpretation of splashback as a genuine feature of the matter distribution, rather than an artefact of galaxy selection or projection effects. Mpetha et al. (2025) used weak lensing profiles from the UNIONS survey to constrain cosmological parameters via the infall region structure of galaxy clusters. Their analysis yielded splashback radii of $1.59^{+0.16}_{-0.13}$, $1.30^{+0.25}_{-0.13}$, and $1.45\pm 0.11\,\mathrm{cMpc}/h$ for three independent cluster samples, and provided constraints of $\Omega_{\mathrm{m}}=0.29\pm 0.05$ and $\sigma_{8}=0.80\pm 0.04$. These results demonstrate that splashback-based lensing analyses can break degeneracies inherent in traditional abundance and shear-based cosmological tests.
However, lensing studies also highlight systematic challenges. Zhang et al. (2023) showed that neglecting the splashback feature in mass profile modelling can bias weak lensing mass estimates, particularly for group-sized halos ($M_{200m}\sim 10^{13.5}\,M_{\odot}$), by more than 0.1 dex. This bias propagates into cosmological analyses, potentially leading to underestimates of $\Omega_{\mathrm{m}}$. Umetsu & Diemer (2017) further demonstrated that the splashback feature is most pronounced when cluster profiles are scaled by $R_{200m}$, and that averaging over physical radii can smear out the steepening signal, underscoring the importance of profile scaling in lensing analyses.
@@
SZ-selected cluster samples offer a redshift-independent, mass-limited catalogue of clusters with minimal projection bias. Studies using clusters from Planck, SPT, and ACT have demonstrated that the splashback radius inferred from the galaxy distribution in SZ-selected samples closely matches expectations from simulations (e.g. Shin et al., 2019; Zürcher & More, 2019; Adhikari et al., 2021). In contrast, optically selected clusters—such as those in redMaPPer samples—often yield smaller splashback radii due to selection and projection effects. Because SZ selection is more tightly correlated with halo mass, it provides an important benchmark for testing the physical interpretation of the splashback feature.
@@
Our analysis of the stellar splashback feature, traced through the total stellar density profiles, adds a new observational window on this boundary. We find that the truncation radius of the stellar orbiting component coincides with that of the dark matter and follows the same dependence on the mass accretion rate, $\Gamma$, in agreement with earlier simulation-based studies (Deason et al., 2021; Diemer, 2025). The steeper fall-off in the stellar profile makes the splashback edge more sharply defined than in lensing profiles, and may be observable through deep imaging of the ICL. These findings validate the use of projected stellar profiles as observational proxies for cluster boundaries. While subject to baryonic effects, the close alignment with dark matter splashback structures highlights the value of stellar tracers for probing cluster dynamics.
@@
Taken together, lensing, SZ, and stellar-based methods provide a powerful set of tools for characterising the outskirts of galaxy clusters. Our results, which show that the stellar splashback radius coincides with that of the dark matter and follows the same accretion rate dependence, suggest that stellar tracers—especially the ICL—can provide a complementary and potentially sharper observational probe of cluster boundaries, provided that projection effects and substructure contamination are properly accounted for.
@@
Each is sensitive to different physical components—total mass, thermal pressure, and stellar material—and affected by distinct systematics. The convergence of splashback measurements across these methods strengthens the interpretation of this feature as a physically meaningful halo boundary and supports its use as a probe of cluster growth and accretion history (e.g. Joshi et al., 2025).

### v2

@@ 4.2 Recovering $r_{\rm t}$ from observable profiles
We calculate the projected density profiles using the same radial binning as we used for the 3D profiles. The angular binning is also done with the same number of bins, but this time in equal angle rather than equal solid angle. We project the entire cluster region along the $x$, $y$ and $z$ axes, and take the mean of the resulting profiles. Since the zoom regions contain only the cluster and its correlated surroundings, there is no contribution to the profile from uncorrelated field galaxies. We therefore do not apply any background subtraction (see Sun et al. 2025 for the effect of the projection window depth on the splashback feature). The calculation of the density slope profiles is the same as for the 3D case. For clarity, we refer to the projected density with the symbol $\Sigma$ and other quantities with upper case, e.g. $R_{\rm t}$ and $B$ for the truncation radius and steepness.
In projection, the shape of the density profile is modified. For the stars, the splashback region and BCG-stellar halo transition are smoothed out. This can be seen in Figure 6, which compares the 3D and projected stellar profile of one of the clusters in the GIZMO_7k run. Despite this, the profiles are still well-fit by Equations 13, generalizing earlier work on the similarity between 3D and projected total profiles (e.g. Merritt et al. 2005). It is reasonable to expect that the parameters of the 3D and projected profiles should be related. Figure 7 compares the truncation radius and truncation steepness from fitting the 3D and projected orbiting profiles (not the total profiles), for the dark matter (left) and stars (right). The projected $R_{\rm t}$ is slightly smaller than the 3D $r_{\rm t}$ by a factor of $\sim 0.92$ for the dark matter and $0.94$ for the stars (black dashed lines). This is similar to the factor of $0.9$ found by Deason et al. (2021) when comparing the 3D and projected steepest slope radii. The projected $B$ of the dark matter is also smaller than its 3D counterpart by a factor of $\sim 0.87$. This may also be the case for the stars, but the scatter is significantly larger.
@@
Figure 8: Comparison between the projected truncation radius obtained by fitting the orbiting profile (labelled component) and the total profile. The points are colored by accretion rate (Equation 4). The cyan points are the median total-profile $R_{\rm t}$ calculated in five equally-spaced bins. The square points are the result of fitting the $\Gamma$-binned median profiles.
Figure 9: $R_{\rm steep}$ and $R_{\rm t}$ as a function of accretion rate (Equation 4). The light blue points are measurements for individual profiles, while the dark blue points are the median and standard deviation in equally-spaced bins in $\Gamma$. The red squares are from fitting the $\Gamma$-binned median profiles. The black dashed line in the top row is the relation from More et al. (2015), rescaled using the factor of $r_{\rm steep}/R_{\rm steep}=0.9$ found by Deason et al. (2021). The black dashed in the bottom row is the same curve further rescaled using the best-fit $r_{\rm t}$$r_{\rm steep}$ relation shown in Figure 4. The black dotted line is the relationship from Diemer (2025), rescaled by the slope of the best-fit line shown in Figure 7.
@@
The $R_{\rm steep}$$\Gamma$ plots feature a small group of points that lie below the relation at low $\Gamma$. These are cases where the $2^{\rm nd}$ caustic—an additional, usually shallower caustic at smaller radius—actually becomes steeper than the $1^{\rm st}$ caustic when viewed in projection. The measurement of $R_{\rm t}$ seems to be unaffected by this, suggesting the fitting method may have an advantage in these cases.
@@ 4.3 The infalling profile
Figure 10: Best-fit parameters for the 3D infalling profiles. The bottom-left (red panels) and top-right (purple panels) show the correlations between each parameter for the dark matter and stars, respectively. The gray diagonal panels compare the parameters between the stars and dark matter. The square points are the result of fitting the median of profiles binned by $\Gamma$, as specified in the legend.
@@
Figure 10 shows the distribution of the parameters from fitting the 3D infalling profiles. The bottom-left (red panels) and top-right (purple panels) show the correlations between each parameter for the dark matter and stars, respectively. The gray diagonal panels compare the parameters between the stars and dark matter. The square points are the result of fitting the median $\Gamma$-binned profiles.
@@ 4.4 The inner profile
In our model, the orbiting density profile of the stars is composed of a Sérsic BCG and a stellar “halo” fit with the truncated Einasto profile of Diemer (2025) (see Section 3.3). As with the unmodified Einasto profile, the shape of the inner profile is controlled by the inner slope exponent $\alpha$ and the scale radius $r_{\rm s}$—the radius at which the logarithmic slope equals 2-2. Unlike for the outer profile parameters, we find that the inner stellar profile parameters do not track those of the dark matter well. The stellar $r_{\rm s}$ tends to underestimate that of the dark matter, while $\alpha$ has significant scatter around the one-to-one line. These comparisons are shown in Figure 13.
At first glance, this discrepancy suggests that the dark matter and stellar profiles may not share a common scale radius, or that a well-defined scale radius may not exist for the stellar component at all. This would not be unexpected given the presence of the central BCG, and the growing importance of dynamical friction and tidal stripping toward the cluster centre, which reshape the stellar and dark matter distributions in different ways.
To test this, we identify $r_{\rm s}$ non-parametrically by locating the peak of the $\rho\,r^{2}$ profile. To do this, we take the median of profiles binned by $\Gamma$, considering both the angular-median profiles used throughout this work and conventional spherically averaged profiles. We further divide the stellar component into stars bound to halos (as identified by ahf) and the remaining diffuse BCG+ICL component. This is shown in Figure 12 for the spherically-averaged case. The stellar profiles, shown in yellow, have been rescaled by the median value of $1.5\times M_{\rm DM}(<R_{200{\rm c}})/M_{\ast}(<R_{200{\rm c}})$ in each bin so that they lie approximately on top of the dark matter curves (black). The density profiles of the stars in satellites have a clear peak that coincides with that of the dark matter, as shown by the logarithmic slope profiles in the bottom panel. Indeed, the satellite profile in the high-accretion-rate bin closely follows the dark matter profile until about $\sim 0.5\,r_{\rm s}$ before falling off more rapidly towards smaller radii dominated by the BCG. In the low-accretion case, $r_{\rm s}/R_{200{\rm m}}$ is smaller (i.e. the concentration is larger) and the fall-off in the satellite density occurs close to $r_{s}$. The BCG also extends to larger $r/R_{200{\rm m}}$, which conceals the peak at these low accretion rates.
These results suggest that, at high accretion rates, the total stellar profile accurately traces the dark matter profile around the scale radius. Meanwhile, at low accretion rates, the BCG encroaches on this region and results in a total profile that is consistently steeper than 2-2. However, the disagreement we find between the stellar and dark matter $r_{\rm s}$ when fitting the orbiting profiles is not a result of this concealing of the peak at low $\Gamma$. Rather, it is likely a consequence of the suppressing of the satellite density around $r_{\rm s}$ by the angular median. This is discussed in more detail in Appendix A.
@@ 5.1 Comparison to Previous Simulations
The coincidence between the $r_{\rm t}$ means that the stellar $r_{\rm t}$$\Gamma$ relation has the same form as that of the dark matter. This makes it a complimentary estimator of the accretion rate in clusters—assuming that $r_{\rm t}$ can be measured accurately. However, in observations, information of the orbiting and infalling components is inaccessible, and $r_{\rm t}$ must instead be estimated by fitting the total profile in projection. Given the freedom allowed by the full nine-parameter model of the total profile, the orbiting and infalling components may not always be accurately recovered.
When fitting the total profile, we find that the projected truncation radius, $R_{\rm t}$, has a standard deviation from the “true” $R_{\rm t}$ (measured from the orbiting profile) of $\sim 0.3\,R_{200{\rm m}}$ for both the dark matter and stars. There may also be a slight $R_{\rm t}$-dependent systematic difference in the case of the stars, which would produce a subtle change in the $R_{\rm t}$$\Gamma$ relation, but this is unlikely to be discernable. While the exponential relationship can still be clearly seen, the scatter is larger than that of the $R_{\rm steep}$$\Gamma$ relationship and so it is likely to be a poorer estimator of the accretion rate in observations.
Figure 12: Median spherically-averaged $\rho r^{2}$ profiles of the dark matter and stars, in two bins of $\Gamma$. The stars have been separated into those contained in satellites—as determined by ahf—and all other stars, i.e. the smooth BCG+ICL. The stellar profiles have been rescaled $1.5\times M_{\rm DM}(<R_{200{\rm c}})/M_{\ast}(<R_{200{\rm c}})$ so that they approximately lie on top of the dark matter curves. The corresponding logarithmic slope profiles are shown in the bottom panel. We emphasise that the profiles that are being averaged here are the spherically-averaged profiles—they have not been binned in angular segments as is done for the other results in this work.
@@ 5.2 Comparison to Observations
Gravitational lensing and measurements of the distribution of cluster galaxies in projection provide complementary approaches to identifying the splashback boundary in galaxy clusters, each offering independent insight into the halo’s outer structure. Weak lensing directly probes the total matter distribution and has revealed a steepening in the density profile consistent with theoretical predictions of the splashback feature (e.g. Chang et al., 2018a; Contigiani et al., 2019) that can be fit with an orbiting-infalling fitting function (e.g. Giocoli et al., 2024; Mpetha et al., 2025). However, lensing measurements are inherently noisy at large radii due to shape noise and projection effects, and often require stacking to obtain robust profiles.
Chang et al. (2018a) conducted one of the earliest joint analyses of galaxy density and weak lensing profiles using DES Year 1 data, detecting a splashback-like steepening in both tracers. They measured a splashback radius of $r_{\mathrm{sp}}=1.34\pm 0.21\,h^{-1}\,\mathrm{Mpc}$, consistent with predictions from simulations and earlier SDSS-based studies. This dual detection reinforced the interpretation of splashback as a genuine feature of the matter distribution, rather than an artefact of galaxy selection or projection effects. Mpetha et al. (2025) used weak lensing profiles from the UNIONS survey to constrain cosmological parameters via the infall region structure of galaxy clusters. Their analysis yielded splashback radii of $1.59^{+0.16}_{-0.13}$, $1.30^{+0.25}_{-0.13}$, and $1.45\pm 0.11\,\mathrm{cMpc}/h$ for three independent cluster samples, and provided constraints of $\Omega_{\mathrm{m}}=0.29\pm 0.05$ and $\sigma_{8}=0.80\pm 0.04$. These results demonstrate that splashback-based lensing analyses can break degeneracies inherent in traditional abundance and shear-based cosmological tests.
Lensing studies also highlight systematic challenges. Zhang et al. (2023) showed that neglecting the splashback feature in mass profile modelling can bias weak lensing mass estimates, particularly for group-sized halos ($M_{200m}\sim 10^{13.5}\,M_{\odot}$), by more than 0.1 dex. This bias propagates into cosmological analyses, potentially leading to underestimates of $\Omega_{\mathrm{m}}$. Umetsu and Diemer (2017) further demonstrated that the splashback feature is most pronounced when cluster profiles are scaled by $R_{200m}$, and that averaging over physical radii can smear out the steepening signal, underscoring the importance of profile scaling in lensing analyses.
@@
Cluster samples selected by their Sunyaev-Zel’dovich (SZ) effect offer a redshift-independent, mass-limited catalogue of clusters with minimal projection bias. Studies using clusters from Planck, SPT, and ACT have demonstrated that the splashback radius inferred from the galaxy distribution in SZ-selected samples closely matches expectations from simulations (e.g. Shin et al., 2019; Zürcher and More, 2019; Adhikari et al., 2021). In contrast, optically selected clusters—such as those in redMaPPer samples—often yield smaller splashback radii due to selection and projection effects (Murata et al., 2020). Because SZ selection is more tightly correlated with halo mass, it provides an important benchmark for testing the physical interpretation of the splashback feature.
@@
Our analysis of the stellar splashback feature, traced through the total orbiting and infalling stellar density profiles, adds a new observational window on this boundary. We find that the truncation radius of the stellar orbiting component coincides with that of the dark matter and follows the same dependence on the mass accretion rate, consistent with earlier simulation-based studies (Deason et al., 2021; Diemer, 2025). The steeper fall-off in the stellar profile makes the splashback edge more sharply defined than in lensing profiles, and may be observable through deep imaging of the ICL.
Observations have not yet reached the surface-brightness limits required to trace the density profiles of galaxy clusters to the splashback radius, though rapid progress is being made. Deason et al. (2021) compared stellar profiles from the C-EAGLE simulations (Barnes et al., 2017) to stacked surface brightness measurements of $z=0.25$ DES Year-1 clusters (Zhang et al., 2019), demonstrating that depths of $\mu\simeq 32$$36~{\rm mag\,arcsec^{-2}}$ are required to probe the splashback regime. These limits are expected to be reached in forthcoming surveys, including LSST (Ivezić et al., 2019).
Given the observational challenges, we have also considered measuring the splashback feature using galaxy number density profiles, which are more readily observable. Because the stellar mass in the outskirts of clusters is dominated by satellites, we expect—and find—very close agreement between the total stellar profile and galaxy number density profiles. We refer the interested reader to Appendix B.
Our findings validate the use of projected stellar profiles as observational proxies for cluster boundaries. While subject to baryonic effects, the close alignment with dark matter splashback structures highlights the value of stellar tracers for probing cluster dynamics.
@@
Taken together, lensing and stellar-based methods provide a powerful set of tools for characterising the outskirts of galaxy clusters. Our results, which show that the stellar splashback radius coincides with that of the dark matter and follows the same accretion rate dependence, suggest that stellar tracers—especially the ICL—can provide a complementary and potentially sharper observational probe of cluster boundaries, provided that projection effects and substructure contamination are properly accounted for.
@@
Each is sensitive to different physical components—total mass, thermal pressure, and stellar material—and affected by distinct systematics. The convergence of splashback measurements across these methods strengthens the interpretation of this feature as a physically meaningful halo boundary and supports its use as a probe of cluster growth and accretion history (e.g. Joshi et al., 2026).